# Compositional Design of Safety Controllers for Large-scale Stochastic Hybrid Systems

### v1

@@ 5. Construction of A-CSBC
      $\leq e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}\big{(}e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\big{)}$
      $=e^{-\gamma_{1_{i}}\tau}\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}}e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}\bar{\mathcal{B}}_{i}(x_{i})+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~{}~{}~{}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
@@
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$ $=e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}^{\prime}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
@@
      $\gamma_{i}$ $=\max\{e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})},e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\},\quad\varphi_{i}=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
@@
      $\eta_{i}$ $=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
@@
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\varphi_{i}\|w_{i}\|^{2}\!\!+\!\eta_{i}.$
@@
      $\leq\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\Big{(}e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\Big{)}$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}\bar{\mathcal{B}}_{i}(x_{i})+\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}\!+\!\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
@@
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$ $=\gamma_{2_{i}}^{\frac{\vartheta_{i}^{\prime}}{\alpha_{2_{i}}}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$
      $=\!\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i}\Big{]}\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
@@
      $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}},\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\},\quad\varphi_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
@@
      $\eta_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
@@
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}+\eta_{i}.$
To complete the proof, we proceed with showing that $10<\gamma_{i}<$ 1 for all Bounds 1-3. One can readily verify that
@@
for Bound 1. In Bound 2, since $10<\alpha_{1_{i}}<$ 1 and $\gamma_{1_{i}}>0$, then we have $e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}<1$. In addition,
@@
Since $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\vartheta_{i}<0$ for any $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$ and by leveraging the continuity property of real numbers, one can always find some $10<\alpha_{1_{i}}<$ 1 such that $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}<0$. Hence, $10<\gamma_{i}<$ 1. Similarly, in Bound 3, one has $\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}<1$ since $\alpha_{2_{i}}>\varepsilon_{2_{i}}$ and $10<\gamma_{2_{i}}<$ 1. Additionally,
@@
6. Computation of CBC
Here, we employ sum-of-squares (SOS) optimization techniques [Par03] and search for a CBC and its corresponding controller for each $\Phi_{i}$. To do so, we assume that $f_{1_{i}}$ and $f_{2_{i}}$ are polynomial functions. The next lemma provides the SOS formulation.

### v2

@@ 5 Construction of A-CSBC
      $\leq e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}\big(e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})$
      $~~~+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\big)$
      $=e^{-\gamma_{1_{i}}\tau}\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}}e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}}\bar{\mathcal{B}}_{i}(x_{i})$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!$
      $~~~+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}(\vartheta_{i}+1)}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $~~~+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $~~~+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $+\!e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}$
      $+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
@@
      $\gamma_{i}$ $=\max\{e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})},e^{-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}}\gamma_{2_{i}}\},$
      $\varphi_{i}$ $=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
@@
      $\eta_{i}$ $=\max\{e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{2_{i}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
@@
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\big|x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!$
      $~~~+\!\varphi_{i}\|w_{i}\|^{2}\!\!+\!\eta_{i}.$
@@
      $\leq\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\Big(e^{-\gamma_{1_{i}}\tau}\bar{\mathcal{B}}_{i}(x_{i})+e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $\quad+e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}\Big)$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}}\bar{\mathcal{B}}_{i}(x_{i})$
      $\quad+\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}\!+\!\gamma_{2_{i}}^{\frac{\vartheta_{i}+1}{\alpha_{2_{i}}}}\!e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}+\|w_{i}\|^{2})$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}$
      $\quad+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}}.$
@@
      $\gamma_{i}$ $=\max\{e^{-\gamma_{1_{i}}\tau}\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}},\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}\},$
      $\varphi_{i}$ $=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\},$
@@
      $\eta_{i}=\max\{\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}$ $\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $+\gamma_{2_{i}}^{\frac{1}{\alpha_{2_{i}}}}e^{-\gamma_{1_{i}}\tau}\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\},$
@@
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big|x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})$
      $~~~+\varphi_{i}\|w_{i}\|^{2}+\eta_{i}.$
To complete the proof, we proceed with showing that $0<\gamma_{i}<1$ for all Bounds 1-3. One can readily verify that
@@
for Bound 1. In Bound 2, since $0<\alpha_{1_{i}}<1$ and $\gamma_{1_{i}}>0$, then we have $e^{-\gamma_{1_{i}}\tau(1-\alpha_{1_{i}})}<1$. In addition,
@@
Since $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\vartheta_{i}<0$ for any $\vartheta_{i}\in\{\varepsilon_{1_{i}},\dots,\varepsilon_{2_{i}}\}$ and by leveraging the continuity property of real numbers, one can always find some $0<\alpha_{1_{i}}<1$ such that $\ln(\gamma_{2_{i}})-\gamma_{1_{i}}\tau\alpha_{1_{i}}\varepsilon_{1_{i}}<0$. Hence, $0<\gamma_{i}<1$. Similarly, in Bound 3, one has $\gamma_{2_{i}}^{\frac{\alpha_{2_{i}}-\varepsilon_{2_{i}}}{\alpha_{2_{i}}}}<1$ since $\alpha_{2_{i}}>\varepsilon_{2_{i}}$ and $0<\gamma_{2_{i}}<1$. Additionally,
@@
Theorem 5.4 establishes that the A-CSBC conditions are fulfilled under the transformation $\mathcal{B}_{i}(x_{i},\vartheta_{i})=\varpi_{i}\bar{\mathcal{B}}_{i}(x_{i})$, provided that condition (5.3) holds. The proposed approach systematically addresses all hybrid transitions—both flows and jumps—by categorizing them into three distinct cases, corresponding to the different decay or growth patterns exhibited by the original CBC.
Remark 5.5.
The introduction of multiple definitions of control barrier certificates in this work serves distinct but complementary purposes within our framework. Definition 3.1 introduces A-CSBCs tailored to the A-SHS model defined in Definition 2.8. While A-CSBCs are necessary to ensure compositional safety using subsystems, their design is complicated by the inclusion of the auxiliary counter $\vartheta_{i}$. In contrast, Definition 5.1 presents a conventional CBC formulation (excluding the effect of $\vartheta_{i}$) for SHS by combining the results of stochastic non-hybrid continuous- and discrete-time systems, which serves as the foundational tool for safety analysis. To bridge the gap between these two definitions, Theorem 5.4 establishes a systematic method to construct A-CSBCs from conventional CBCs by introducing an auxiliary parameter $\varpi_{i}$.
Remark 5.6.
The core challenge in analyzing networked stochastic hybrid systems lies not only in their high dimensionality but also in the hybrid coupling between subsystems, captured by the adversarial inputs $w_{i}$ in (2.4). This coupling significantly impacts the system’s behavior and should be explicitly considered in our formal safety analysis. To this end, our framework incorporates $w_{i}(t)$ throughout the analysis: they are integrated in our development in Lemma 5.3, their bounds are formally established in Theorem 4.2, and their impact is captured by the parameter $\varphi_{i}$, introduced in Definition 3.1. This parameter is then properly designed in Theorem 5.4—as shown in its proof—under three distinct scenarios (Bounds 1–3), that can cover the hybrid nature of dynamics.
6 Computation of CBC
Here, we employ sum-of-squares (SOS) optimization techniques (Parrilo, 2003) and search for a CBC and its corresponding controller for each $\Phi_{i}$. To do so, we assume that $f_{1_{i}}$ and $f_{2_{i}}$ are polynomial functions. The next lemma provides the SOS formulation.